# Asynchronous Large Language Model Enhanced Planner for Autonomous Driving

### v2

@@ 5.1.2 Implementation Details
Regarding implementation details, all experiments were conducted in the closed-loop reactive setting, where agents in scenarios could be equipped with an IDM Treiber et al. (2002) Planner, enabling them to react to ego vehicle’s maneuvers. The simulation frequency is $10Hz$, at each iteration, the predicted trajectory has a time horizon of $8s$. We follow the closed-loop metrics proposed in nuPlan challenge, which is detailed in the supplementary material. For model settings, our AsyncDriver is built based on Llama2-13B Touvron et al. (2023), LoRA Hu et al. (2021) was configured with $R=8$ and $\alpha=32$. We use the AdamW optimizer and warm-up with decay scheduler with learning rate $0.0001$.
(a) score variation with intervals
(b) average inference time variation with intervals
@@ 5.2.1 Hard20 Evaluation
As illustrated in Tab. 1, our approach AsyncDriver achieved the highest performance on Hard20 compared with existing planners, leading to an improvement of $4.6\%$ over GameFormer Huang et al. (2023) about $2.95$ in score, even surpassing the current SOTA rule-based planners. For fair comparison, given the significant impact of trajectory refinement and alignment in closed-loop evaluations, we adapt the PDM Dauner et al. (2023) scorer to our AsyncDriver (denoted as AsyncDriver*), which lead to an improvement of $5.3\%$ and $5.1\%$ over PDM-Hybrid Dauner et al. (2023) and PDM-Closed Dauner et al. (2023) respectively, equivalent to approximately $3.43$ and $3.30$ in score, and an $8.7\%$ increase (approximately $5.43$ in score) over the learning-based planner GameFormer. From a different perspective, Tab. 2 illustrates the scores of each individual scenario type on the hard20 split, as well as a comparison with existing planners. It is evident that our solution has delivered exceptional results in the majority of scenario types.
The quantitative results demonstrate that the high-level feature extracted by our Scene-Associated Instruction Feature Extraction Module effectively enhances the performance of real-time planners in closed-loop evaluation. Moreover, the detailed metric scores reveal that our approach significantly improves in drivable area performance compared to GameFormer, with a $3.23$ point increase, which proves that our model’s superior capability in accurately identifying and navigating viable driving spaces, attributing to its advanced scene contextual understanding abilities. Similarly, compared to PDM, AsyncDriver* exhibits remarkable superiority in Time to Collision (TTC) metrics, outperforming it by $4.39\%$, approximately $3.23$ in score, This indicates the model’s enhanced predictive accuracy, which is crucial for ensuring safer driving by effectively forecasting and reducing the potential collision scenarios.
@@ 5.2.3 Instruction Following
Fig. 4 shows the reactions of our method to different routing instructions, demonstrating its capability in instruction following. Fig. 4a illustrates the outcomes predicted when the scene employs conventional routing instructions. We note that the ego vehicle decelerates slightly, a maneuver that reflects the common sense of slowing down for curves. Nonetheless, given the open road conditions, the ego vehicle sustains a comparatively high velocity. In contrast, Fig. 4b depicts the scenario where stop serves as the routing instruction. Remarkably, even without the presence of external obstacles, the ego vehicle promptly executes a braking response to the command, reducing its speed from $10.65m/s$ to $1.06m/s$ within a mere $6$ seconds. Consequently, it is evident that our AsyncDriver can function as a linguistic interaction interface, providing the ability of precisely interpreting and following human instructions to circumvent anomalous conditions.
(a) receiving conventional routing instructions
(b) receiving stop instruction
@@ 5.3 Ablation Study
In this section, we investigate the effectiveness of individual components in AsyncDriver. An attempt of integrating a simple MLP as prediction head subsequent to LLM for the planning task resulted in a significant degradation of closed-loop performance. This indicates that taking simple trajectory regression task as supervision cannot correctly align multi-modal information, making it difficult to tap into LLM’s inherent knowledge and its potential for scene interpretation and reasoning. Subsequently, we removed the MLP structure in favor of a real-time planner and progressively added four following structures: (i) Adaptive Injection Block, (ii) Alignment Assistance Module, (iii) LoRA, and (iv) pre-trained LoRA. As delineated in Tab. 3, each introduced module was instrumental in augmenting performance. Notably, the Alignment Assistance Module and the pre-trained LoRA weights, yielded performance uplifts of $0.94$ and $0.97$ scores respectively, have the most significant contributions.
Table 3: Component ablation study. MLP, RT-Planner, Ada, Align, LoRA, and LoRAPre represent i) direct regression of waypoints using MLP, ii) integration of a real-time planner, iii) Adaptive Injection Block, iv) Alignment Assistance Module, v) incorporation of LoRA, and vi) pre-trained LoRA, respectively.
@@ 6 Conclusions
Broader Impacts And Limitations
@@
In this supplementary material, we first present the templates for QAs and prompts of pre-training and fine-tuning data (Appendix A). Then, we provide detailed metric information for our closed-loop evaluation (Appendix B), as well as the specific 14 scenario types defined in nuPlan challenges Motional (2023) (Appendix C). In the end, we report additional quantitative results conducted to further substantiate the efficacy of our approach (Appendix D).
Appendix A Data Templates
A.1 Pre-training Data
A.1.1 Planning-QA
We assessed the capabilities of Llama2-13B Touvron et al. (2023) via a zero-shot methodology and discovered that its extensive pre-training data provides a solid foundation in traffic rule comprehension. However, its limited mathematical prowess poses challenges in grasping and deducing the connections between instructions and numerical expressions. To address this, we introduced a language-based Planning-QA dataset aimed at transitioning LLM from a general model to a specialized model adept in autonomous driving planning. This enhancement focuses on refining its capability in instruction interpretation and reasoning.
Concretely, we delineated the level of autonomous driving planning into three granularities: 1) high-level instructions: formulated through velocity commands including stop, accelerate, decelerate, maintain speed, and routing commands including turn left, turn right, go straight, 2) control: assessing the values of velocity and acceleration, 3) and waypoint: encompassing a series of points. Six question types were devised to articulate the transitional relationships across the high-level instructions - control - waypoint spectrum, and each QA-pair is adapted based on log data from nuPlan Caesar et al. (2021). Fig. S1a illustrates the universal system prompt template applicable to all questions, whereas Fig. S1b-S2b display specific examples for each question type alongside their respective answer, substituting <Question> and <Answer> within the system prompt.
(a) system prompt template for Planning-QA dataset
(b) waypoints - high-level instructions
(c) waypoints - control
(d) high-level instructions - waypoint
(e) high-level instructions - control
Figure S1: Question types
(a) control - waypoint
(b) control - high-level instructions
@@ A.1.1 Planning-QA
A.1.2 Reasoning1K
Additionally, we have crafted 1,000 QA-pairs using GPT-4, tailored to tasks that require not just decision-making but also in-depth reasoning. Each sample includes a question that sets the context, provides scene descriptions, outlines considerations based on traffic rules, details the anchor trajectory from real-time planners, and specifies requirements for the answers, while the answer delivers trajectory predictions along with the decisions and the underlying thought process. By training with mixed of Reasoning1K and Planning-QA, we have substantially enriched the dataset’s diversity and complexity. This approach aims to bolster LLM’s ability to deduce answers through logical reasoning, seamlessly integrating traffic rules with detailed, scenario-based analysis. Fig. S3a and Fig. S3b showcase an example of the input and its corresponding target from the dataset, respectively.
(a) example of input for Reasoning1K
(b) example of corresponding target for Reasoning1K
Figure S3: An example from Reasoning1K, where S3a displays the input, and S3b shows the corresponding target. The reasoning-related parts are highlighted in color, with the same color in both input and target representing the cause and effect in the reasoning process respectively.
A.2 Fine-tuning Data
@@
Appendix B Closed-loop Metrics
In this section, we elucidate the metrics of closed-loop evaluation deployed within the nuPlan Caesar et al. (2021) framework. The calculation for the composite metric Score is detailed in Formula 6.
@@ Appendix B Closed-loop Metrics
Appendix C Scenario Types
In this section, we delineate $14$ scenario types employed in both the training and evaluation which is consistent with the nuPlan Challenge 2023 Motional (2023). These scenarios encompass a broad spectrum of driving conditions to evaluate the adaptability and planning capabilities of autonomous vehicles. The scenarios are categorized as follows:
@@ Appendix C Scenario Types
Appendix D Visualization
Fig. S5 additionally showcases a series of our visualization results and demonstrates its superior performance to GameFormer Huang et al. (2023) across multiple metrics.
(a) GameFormer collides with other vehicles at 6.5s; AsyncDriver first yield to the vehicle going straight, then passes smoothly.
(b) GameFormer, upon reaching 11.5s, finds it difficult to evade pedestrians ahead, resulting in a collision at 15.0s; AsyncDriver, anticipating the situation in advance, passes smoothly before the pedestrians approach.
(c) GameFormer makes a misjudgment about the traffic light status at 5.0s, noticeably decelerating, leading to its failure to enter the intersection at 12.5s due to slow progress; AsyncDriver maintains a steady speed and enters the intersection smoothly.
(d) GameFormer, after the traffic light turns from red to green at 11.0s, fails to start promptly and remains stationary at 15.0s; AsyncDriver starts on time and enters the intersection.
(e) At 12.5 seconds, when a pedestrian appears ahead, GameFormer does not evade or decelerate; AsyncDriver clearly brakes.
(f) GameFormer gets too close to the adjacent car while turning, increasing the risk of collision; AsyncDriver travels almost parallel to the neighboring car, greatly enhancing safety.
(g) GameFormer turns out of the drivable area; AsyncDriver does not.
(h) GameFormer exceeds the drivable area; AsyncDriver does not.

### v3

@@ 5.1.2 Implementation Details
Regarding implementation details, all experiments were conducted in the closed-loop reactive setting, where agents in scenarios could be equipped with an IDM [39] Planner, enabling them to react to ego vehicle’s maneuvers. The simulation frequency is $10Hz$, at each iteration, the predicted trajectory has a time horizon of $8s$. We follow the closed-loop metrics proposed in nuPlan challenge, which is detailed in the supplementary material. For model settings, our AsyncDriver is built based on Llama2-13B [38], LoRA [15] was configured with $R=8$ and $\alpha=32$. We use the AdamW optimizer and warm-up with decay scheduler with learning rate $0.0001$.
@@ 5.2.1 Hard20 Evaluation
As illustrated in Tab. 1, our approach AsyncDriver achieved the highest performance on Hard20 compared with existing planners, leading to an improvement of $4.6\%$ over GameFormer [18] about $2.95$ in score, even surpassing the current SOTA rule-based planners. For fair comparison, given the significant impact of trajectory refinement and alignment in closed-loop evaluations, we adapt the PDM [10] scorer to our AsyncDriver (denoted as AsyncDriver*), which lead to an improvement of $5.3\%$ and $5.1\%$ over PDM-Hybrid [10] and PDM-Closed [10] respectively, equivalent to approximately $3.43$ and $3.30$ in score, and an $8.7\%$ increase (approximately $5.43$ in score) over the learning-based planner GameFormer. From a different perspective, Tab. 2 illustrates the scores of each individual scenario type on the hard20 split, as well as a comparison with existing planners. It is evident that our solution has delivered exceptional results in the majority of scenario types.
Quantitative results show that the high-level features extracted by our Scene-Associated Instruction Feature Extraction Module significantly enhance real-time planners’ performance in closed-loop evaluation. Detailed metrics reveal that our approach improves drivable area performance by $3.23$ points compared to GameFormer, demonstrating superior capability in identifying and navigating viable driving spaces due to advanced scene contextual understanding. Additionally, AsyncDriver* outperforms PDM in Time to Collision (TTC) metrics by $4.39$%, approximately $3.23$ points, indicating enhanced predictive accuracy, which is crucial for ensuring safer driving by effectively forecasting and reducing the potential collision scenarios.
@@ 5.2.3 Instruction Following
Fig. 4 shows the reactions of our method to different routing instructions, demonstrating its capability in instruction following. Fig. 4(a) illustrates the outcomes predicted when the scene employs conventional routing instructions. We note that the ego vehicle decelerates slightly, a maneuver that reflects the common sense of slowing down for curves. Nonetheless, given the open road conditions, the ego vehicle sustains a comparatively high velocity. In contrast, Fig. 4(b) depicts the scenario where stop serves as the routing instruction. Remarkably, even without the presence of external obstacles, the ego vehicle promptly executes a braking response to the command, reducing its speed from $10.65m/s$ to $1.06m/s$ within a mere $6$ seconds. Consequently, it is evident that our AsyncDriver can function as a linguistic interaction interface, providing the ability of precisely interpreting and following human instructions to circumvent anomalous conditions.
@@
Table 3: Replacing LLM with non-pretrained cross attention. Utilize map information and learnable instruction embeddings to perform cross-attention, replacing the features generated by the LLM in the Scene-Associated Instruction Feature Extraction Module.
Hidden Feature Dimensions Instruction Embedding LLM Score
256 63.27
5120 63.59
5120 65.00
@@ 5.3 Ablation Study
In this section, we explore the necessity of LLM and investigate the effectiveness of individual components in AsyncDriver.
5.3.1 Necessity of LLM
We conducted experiments by replacing the LLM with transformer blocks of different dimensions (256 and 5120), incorporating learnable routing instruction embeddings. Meanwhile, the Alignment Assistance Module and Adaptive Injection Block remained unchanged. The results, shown in Tab. 3, indicate that despite a 20-fold increase in the number of transformer parameters, the performance only marginally improved from 63.27 to 63.59. In contrast, our AsyncDriver achieved a performance of 65.00, highlighting the significant impact of pretrained knowledge from LLM.
5.3.2 Ablation of Components
Integrating a simple MLP as a prediction head after the LLM for planning significantly degraded the performance, indicating that simple trajectory regression cannot effectively align multi-modal information, thus failing to leverage the LLM’s knowledge for scene reasoning. We replaced the MLP with a real-time planner and progressively added four structures: (i) Adaptive Injection Block, (ii) Alignment Assistance Module, (iii) LoRA, and (iv) pre-trained LoRA. As shown in Tab. 4, each module improved performance, with the Alignment Assistance Module and pre-trained LoRA weights contributing the most, yielding score increases of $0.94$ and $0.97$, respectively.
Table 4: Component ablation study. MLP, RT-Planner, Ada, Align, LoRA, and LoRAPre represent i) direct regression of waypoints using MLP, ii) integration of a real-time planner, iii) Adaptive Injection Block, iv) Alignment Assistance Module, v) incorporation of LoRA, and vi) pre-trained LoRA with Reasoning1K, respectively.
@@ 6 Conclusions
6.0.1 Broader Impacts And Limitations.
@@
Acknowledgements
This research is supported in part by the National Science and Technology Major Project (No. 2022ZD0115502), the National Natural Science Foundation of China (No. 62122010, U23B2010), Zhejiang Provincial Natural Science Foundation of China (No. LDT23F02022F02), Beijing Natural Science Foundation (No. L231011), Beihang World TOP University Cooperation Program, and Lenovo Research.
@@ Acknowledgements
In this supplementary material, we first present the templates for QAs and prompts of pre-training and fine-tuning data (Appendix 0.A). Then, we provide detailed metric information for our closed-loop evaluation (Appendix 0.B), as well as the specific 14 scenario types defined in nuPlan challenges [30] (Appendix 0.C). In the end, we report additional quantitative results conducted to further substantiate the efficacy of our approach (Appendix 0.D).
Appendix 0.A Data Templates
0.A.1 Pre-training Data
0.A.1.1 Planning-QA
We assessed the capabilities of Llama2-13B [38] via a zero-shot methodology and discovered that its extensive pre-training data provides a solid foundation in traffic rule comprehension. However, its limited mathematical prowess poses challenges in grasping and deducing the connections between instructions and numerical expressions. To address this, we introduced a language-based Planning-QA dataset aimed at transitioning LLM from a general model to a specialized model adept in autonomous driving planning. This enhancement focuses on refining its capability in instruction interpretation and reasoning.
Concretely, we delineated the level of autonomous driving planning into three granularities: 1) high-level instructions: formulated through velocity commands including stop, accelerate, decelerate, maintain speed, and routing commands including turn left, turn right, go straight, 2) control: assessing the values of velocity and acceleration, 3) and waypoint: encompassing a series of points. Six question types were devised to articulate the transitional relationships across the high-level instructions - control - waypoint spectrum, and each QA-pair is adapted based on log data from nuPlan [4]. Fig. 1(a) illustrates the universal system prompt template applicable to all questions, whereas Fig. 2(a)-3(a) display specific examples for each question type alongside their respective answer, substituting <Question> and <Answer> within the system prompt.
@@ 0.A.1.1 Planning-QA
0.A.1.2 Reasoning1K
Additionally, we have crafted 1,000 QA-pairs using GPT-4, tailored to tasks that require not just decision-making but also in-depth reasoning. Each sample includes a question that sets the context, provides scene descriptions, outlines considerations based on traffic rules, details the anchor trajectory from real-time planners, and specifies requirements for the answers, while the answer delivers trajectory predictions along with the decisions and the underlying thought process. By training with mixed of Reasoning1K and Planning-QA, we have substantially enriched the dataset’s diversity and complexity. This approach aims to bolster LLM’s ability to deduce answers through logical reasoning, seamlessly integrating traffic rules with detailed, scenario-based analysis. Fig. 4(a) and Fig. 4(b) showcase an example of the input and its corresponding target from the dataset, respectively.
Figure S4: An example from Reasoning1K, where 4(a) displays the input, and 4(b) shows the corresponding target. The reasoning-related parts are highlighted in color, with the same color in both input and target representing the cause and effect in the reasoning process respectively.
0.A.2 Fine-tuning Data
@@
Appendix 0.B Closed-loop Metrics
In this section, we elucidate the metrics of closed-loop evaluation deployed within the nuPlan [4] framework. The calculation for the composite metric Score is detailed in Formula 6.
@@ Appendix 0.B Closed-loop Metrics
Appendix 0.C Scenario Types
In this section, we delineate $14$ scenario types employed in both the training and evaluation which is consistent with the nuPlan Challenge 2023 [30]. These scenarios encompass a broad spectrum of driving conditions to evaluate the adaptability and planning capabilities of autonomous vehicles. The scenarios are categorized as follows:
@@ Appendix 0.C Scenario Types
Appendix 0.D Visualization
Fig. S5 additionally showcases a series of our visualization results and demonstrates its superior performance to GameFormer [18] across multiple metrics.